OpenStudy (anonymous):

Find the present value of the deposit. $12,000 at 14% compounded quarterly for 8 years. The formula we are given is A = P [1+(r/m)]^(tm)

OpenStudy (anonymous):

\[12000(1+\frac{.14}{4})^{8\times 4}\]

OpenStudy (anonymous):

ok what is up with latex?

OpenStudy (anonymous):

\[x^2\]

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

lemme try another browser

OpenStudy (anonymous):

\[x^2\]

OpenStudy (anonymous):

oh seems to work now. so first answer is the right one

OpenStudy (anonymous):

use a calculator and judicious use of parentheses to compute \[12000(1+\frac{.14}{4})^{32}\]

OpenStudy (anonymous):

I just figured out what I was doing wrong. Seems as if you get a different answer when you put the exponent as 8(4) as 32. Thanks for your help.

OpenStudy (anonymous):

yes because unless you use parentheses it is going to raise to the power of 8 and then multiply the result by 4